HETAV PATEL

Grade 11,

A charge Q is divided into two parts. The two charges kept at a distance apart have a maximum coulombian repulsion. Then the ratio of Q and one of the parts is given by (a) 1 : 2 (b) 2 : 1 (c) 1 : 4 (d) 4 : 1


A charge Q is divided into two parts. The two charges kept at a distance apart have a maximum coulombian


repulsion. Then the ratio of Q and one of the parts is given by


(a) 1 : 2 (b) 2 : 1 (c) 1 : 4 (d) 4 : 1


Grade:11

1 Answers

Anirban Mukherjee
11 years ago

F=q(Q-q)/x^2

dF/dq =0

Q-2q=0

q=Q/2